OpenStudy (anonymous):

Jesse has constructed a huge cylindrical can with a diameter of 60 ft. The can is being filled with water at a rate of 450 ft3/min. How fast is the depth of the water increasing? (Hint: The volume of water in the cylinder is determined by πr2h where r is the radius and h is the depth of the water )

OpenStudy (anonymous):

450 ft^3/min=surface area

OpenStudy (anonymous):

surface area=pir^2

OpenStudy (campbell_st):

this is a related rates question. so you know the the radius is 30ft then the volume is \[V = \pi r^2h\] so you need the derivative of the volume with respect to height. or \[\frac{dV}{dh}\] then take the reciprocal to get \[\frac{dh}{dV}\] and you know \[\frac{dV}{dt} = 450 \] then you need the chain rule \[\frac{dh}{dt} = \frac{dh}{dV} \times \frac{dV}{dt}\] you can plug the radius of 30 in at any time. hope it helps

OpenStudy (anonymous):

Was für ein Mensch bist du? Bist du Lehrer? Denn das ist so beeindruckend .